(π‘₯ βˆ’ 𝑦)𝑦2𝑝 + (𝑦 βˆ’ π‘₯)π‘₯2π‘ž = (π‘₯2 + 𝑦2)𝑧,

Find the Integra surface when π‘₯𝑧 = π‘Ž3 (π‘Ž > 0), 𝑦 = 0.



1
Expert's answer
2021-05-31T16:39:21-0400

"\frac{dx}{x-y)y^2}=\frac{dy}{(y-x)x^2}=\frac{dz}{(x^2+y^2)z}"

"\frac{dx}{x-y)y^2}=\frac{dy}{(y-x)x^2}"

"x^2dx=-y^2dy"

"x^3+y^3=C_1"

"\frac{dx-dy}{xy^2-y^3-yx^2+x^3}=\frac{dz}{(x^2+y^2)z}"

"\frac{d(x-y)}{(x-y)(x^2+y^2)}=\frac{dz}{(x^2+y^2)z}"

"\frac{d(x-y)}{x-y}=\frac{dz}{z}"

"ln(x-y)=ln(z)+ln(C_2)"

"C_2=\frac{x-y}{z}"

"F(x^3+y^3,\frac{x-y}{z})=0"

For xz=a3, y=0

"x^3=C_1"

"x^2=a^3C_2"

The integral surfaceΒ 

x=C